peter started driving at 4:50 and is driving at an average speed of 30 mi/hr. dina is driving at an average of 40 mi/ hr. and she started driving at 5:20 if they both started driving from the same point, after how many hours would the meet?

Answer by ankor@dixie-net.com(12701)

peter started driving at 4:50 and is driving at an average speed of 30 mi/hr. dina is driving at an average of 40 mi/ hr. and she started driving at 5:20 if they both started driving from the same point, after how many hours would the meet?
:
From the information given we know that d's travel time was a half hour less than p's travel time
:
Let t = p's travel time
then
(t-.5) = d's travel time
:
We d catches up with p, the will have traveled the same distance
write a distance equation from this fact: Dist = speed * time
:
d's distance = p's dist
40(t-.5) = 30t
40t - 20 = 30t
40t - 30t = 20
t = 20/10
t = 2 hrs is p's travel time;
:
d catches up with p at 6:50
:
Check solution by confirming that the distances are equal
30(2) = 60 mi
40(1.5) = 60 also